How do you solve a problem like this? y= 3x and 3x+4y= 30
Answer by neatmath(302)

The easiest way is by substitution.

y=3x

3x+4y=30

Since we know what y equals in the first equation (3x), we will substitute that in for the value of y in the second equation.

So we now have

3x+4*(3x)=30

3x+12x=30

15x=30

x=2

Since we know what x has to be, we can just substitute that value into either equation to find the value of y.

y=3x

y=3*2

y=6

The solution to this system of equations is x=2 and y=6, or the point (6,2)

When dealing with a system of 2 linear equations, the solution to the system will always be a singe point, unless the two lines are parallel or are actually the same line. :)

I hope this helps!